2x2+3x-5=0
I'm supposed to solve this equation by using the quadratic formula the 2x2 is two x squared. thanks sarah

Answer by fractalier(2101)

Okay from
2x^2 + 3x - 5 = 0
we can factor this into
(x - 1)(2x + 5) = 0
and so
x = 1, -5/2